The annual cost in dollars for removing p% of the toxic chemicals from a towns water supply is given by the formula
C(P)=500,000/100-P
A. Use the formula to find C(99.5) and C(99.9).

Answer by checkley77(12569)

C(P)=500,000/100-P
C(99.5)=500,000/(100-99.5)
C(99.5)=500,000/.5
C(99.5)=$1,000,000 ANS.
C(99.9)=500,000/9100-99.9)
C(99.9)=500,000/.1
C(99.9)=$5,000,000 ANS.